OpenStudy (anonymous):

20x-25/3b-4 / 4a^2x-5a^2/16-9b^2

OpenStudy (anonymous):

O.K.

OpenStudy (anonymous):

(20x-25)/(3b-4) / (4a^2x-5a^2)/(16-9b^)

hero (hero):

Good job

hero (hero):

So we have \[\frac{20x - 25}{3b - 4} \div\ \frac{4a^2x - 5a^2}{16-9b^2}\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

(5(4x-5)

hero (hero):

Use the reciprocal property to re-write the expression as \[\frac{20x - 25}{3b - 4} \times\ \frac{16-9b^2 }{4a^2x - 5a^2}\]

OpenStudy (anonymous):

Without negative exponents ?

hero (hero):

If there are no negative exponents in the original expression, then we shouldn't have to worry about them. Usually, the goal is to try to avoid them or get rid of them.

hero (hero):

We don't seem to be dealing with any at the moment. It seems that here, the only thing we have to concentrate on is simplifying the expression.

OpenStudy (anonymous):

It becomes (5(4x-5)/(3b-4) / (a^2(4x-5)/((4-3b)*(4+3b))

OpenStudy (anonymous):

Then we rewrite it, like you said.

hero (hero):

For convenience, let's re-write it as \[\frac{(4 - 3b)(4 + 3b)}{(3b - 4)} \times \frac{5(4x - 5)}{a^2(4x - 5)}\]

hero (hero):

Also, let us take out a negative in the denominator so that we can re-write it as: \[\frac{(4 - 3b)(4 + 3b)}{-(4 - 3b)} \times \frac{5(4x - 5)}{a^2(4x - 5)}\] Then: \[-\frac{(4 - 3b)(4 + 3b)}{(4 - 3b)} \times \frac{5(4x - 5)}{a^2(4x - 5)}\] This way it is easier to see what cancels: \[-\frac{\cancel{(4 - 3b)}(4 + 3b)}{\cancel{(4 - 3b)}} \times \frac{5\cancel{(4x - 5)}}{a^2\cancel{(4x - 5)}}\]

OpenStudy (anonymous):

Thank you.

hero (hero):

And thus the result is obvious from here: \[-\frac{5(4 + 3b)}{a^2}\]

hero (hero):

You're welcome